# A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene

### v1

@@ A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene
Figure 1: Multi-dimensional phase space of rhombohedral hexalayer graphene. (a) Schematic of the R6G heterostructure. The red arrow indicates the out-of-plane magnetic field $B_{\perp}$. (b) Schematic phase diagram defined by carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. (c, d) $n$$D$ maps of longitudinal resistance measured at $B=0$, with current flowing along the principal axes of (c) minimum and (d) maximum conductivity. (e) Annotated $n$$D$ map highlighting prominent phases corresponding to panels (c, d). (f, g) $n$$B$ maps measured at $D=967$ mV/nm, showing (f) longitudinal resistance $R_{xx}$ and (g) transverse resistance $R_{xy}$. White dashed lines trace the trajectories of integer quantum Hall states, exhibiting integer Streda slopes. The boundaries of regime I, highlighted by the black dashed lines, follow a Streda slope of $t=2.5$. In panel (f), white dotted lines show Streda slopes of 2 and 3, which clearly deviate from the boundaries of regime I.
@@
Figure 2: $B_{\perp}$-stabilized superconductivity and coexisting quantum Hall state. $n$$D$ maps of (a) $R_{xx}$ and (b) $R_{xy}$ measured at different values of $B_{\perp}$. The dashed line outlines the boundary of regime I, while the dash-dotted line encloses the superconducting phase. (c) Line cuts of $|R_{xy}|$ as a function of $n$ at $D=1000$ mV/nm, measured at various values of $B_{\perp}$. (d, e) Temperature dependence of $R_{xx}$ (red) and $R_{xy}$ (blue) measured from (d) the $B_{\perp}$-stabilized superconducting phase and (e) the adjacent quantum Hall state.
@@
Figure 3: A sequence of half-integer quantum Hall states. (a, b) $n$$B$ map of $R_{xx}$ (top panel) and $R_{xy}$ (bottom panel) measured at (a) $D=980$ mV/nm and (b) $D=1000$ mV/nm. The superconducting region is enclosed by a white dashed line in the top panel, while the sequence of quantum Hall states forms a Landau fan, marked by black dashed lines. (c, d) Temperature dependence of $R_{xy}$ measured at different $B_{\perp}$ values within (c) the quantum Hall states at both integer and half-integer Landau level fillings and (d) the superconducting state.
[…]
@@
When superconductivity persists to high magnetic fields, it opens the door to a rare and intriguing regime in which superconductivity coexists with another hallmark of two-dimensional (2D) electron systems: the quantum Hall effect. The interplay between these two phenomena has long been proposed as a key platform for realizing fault-tolerant quantum computation, owing to the potential emergence of non-Abelian anyons and topological superconductivity Nayak et al. (2008); Das Sarma et al. (2015); Peralta Gavensky et al. (2020); Alicea and Fendley (2016). Over the past decade, this coexistence has been explored extensively in engineered hybrid systems, where superconducting and quantum Hall regions are coupled across artificial interfaces Amet et al. (2016); Lee et al. (2017); Gül et al. (2022); Zhao et al. (2023), as well as in 2D material heterostructures hosting superconductivity and topological order at zero magnetic field Stepanov et al. (2020); Choi et al. (2025). Despite these advances, achieving an intrinsic intertwinement of superconductivity and the quantum Hall effect within a single 2D electronic system—without relying on interface engineering—remains a central open challenge. A recent experimental report has identified such coexistence near zero magnetic field, sparking renewed interest in the possibility of anyonic superconductivity and its topological implications  Xu et al. (2025); Nosov et al. (2025); Pichler et al. (2025).
@@
Facilitated by a dual-encapsulated heterostructure, as illustrated in Fig. 1a, the low-temperature phase space of R6G is governed by three independently tunable parameters: carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. Superconductivity and the sequence of quantum Hall states are both confined within a unique boundary, which encircles a tilted cylindrical volume within this $n$$D$$B$ space, as illustrated in Fig. 1b. For simplicity, we will refer to this portion of the phase space as regime I.
@@
We begin by defining the boundary of regime I through its distinct transport signatures. To this end, we perform measurements across multiple two-dimensional slices of the $n$$D$$B$ parameter space. In particular, we examine $n$$D$ planes at five different values of out-of-plane magnetic field $B_{\perp}$ (blue dashed lines in Fig. 1b), as well as $n$$B$ planes at three fixed values of displacement field $D$ (red dashed lines in Fig. 1b).
@@
At zero magnetic field ($B_{\perp}=0$), the cross section of regime I is bounded by the electronic smectic phase and the striped superconducting state, both previously identified through angle-resolved transport studies Morissette et al. (2025a). The transport response across this regime is characterized in Figs. 1c–d, which show longitudinal resistance measured across the $n$$D$ plane at $B_{\perp}=0$, along two orthogonal principal axes corresponding to maximum and minimum conductivity. The smectic phase occupies the region of pronounced transport anisotropy (see Fig. M2b), highlighted by the red-shaded area in Fig. 1e. Within this anisotropic regime, two superconducting phases—denoted SC i and SC ii—emerge concurrently with the smectic order and inherit its strong directional character  Morissette et al. (2025a).
@@
The $B_{\perp}=0$ cross-section of regime I appears as a metallic phase situated between SC i and SC ii, enclosed by black dashed lines in Figs. 1c–e. As $B_{\perp}$ increases, the boundaries of regime I trace a well-defined trajectory in the $n$$B$ plane, marked by peaks in the longitudinal resistance $R_{xx}$ and abrupt jumps in the Hall resistance $R_{xy}$. These boundaries, indicated by black dashed lines in Fig. 1f-g, exhibit a Streda slope of $t=2.5$  Středa and Smrčka (1983); Widom (1983); MacDonald and Středa (1984).
@@
Within the stripe regime, two additional metallic phases—regime II and regime III—are highlighted in Figs. 1c–e. These phases resemble regime I in their transport characteristics and exhibit well-defined boundaries in the $n$$B$ plane, following distinct Streda slopes: $t=2$ for regime II and $t=3.5$ for regime III (see Fig. M6 for further discussion). The presence of such sharply delineated, evolving boundaries is intriguing. Although a well-defined Streda slope often signals a topological phase, we do not observe transport features typically associated with incompressible quantum Hall states Huang et al. (2025). We revisit this issue later in the context of additional experimental evidence.
@@
Figures 2a–b show $n$$D$ maps of the longitudinal and Hall resistances measured at various values of $B_{\perp}$. At each field, regime I is delineated by peaks in $R_{xx}$, defining a boundary that shifts systematically with $B_{\perp}$ in both carrier density $n$ and displacement field $D$. These evolving boundaries are marked by dashed lines in Figs. 2a–b. Most notably, upon applying a non-zero $B_{\perp}$, a superconducting pocket emerges on the low-density side of regime I, while a sequence of quantum Hall states appears on the high-density side.
The superconducting phase—enclosed by dash-dotted lines—is identified by vanishing longitudinal resistance $R_{xx}$ and zero Hall resistance $R_{xy}$, appearing as dark blue and white, respectively, in the color maps. Figure 2c presents line cuts of the transport response as a function of $n$, taken at fixed $D=1000$ mV/nm and varying $B_{\perp}$. At $B_{\perp}=0$, the system remains resistive across the entire density range. In contrast, at $B_{\perp}=2.5$ and 3.33.33.3 T, a superconducting dome is clearly visible, characterized by a sharp drop to zero resistance. Upon further increasing $B_{\perp}$ to 555 T, the superconducting phase is fully suppressed.
@@
Remarkably, both the superconducting and quantum Hall phases exhibit sharply defined onsets as a function of temperature (Fig. 2d–e). The $R$$T$ curves in Fig. 2d display behavior consistent with a superconducting transition, characterized by a critical temperature $T_{c}\approx 0.5$ K. In contrast, the similarly sharp onset observed for the quantum Hall state in Fig. 2e is highly unusual. Conventional quantum Hall states typically display a gradual thermal activation in $R_{xx}$, while $R_{xy}$ remains largely quantized and temperature-independent (see Fig. M5). The abrupt temperature dependence observed here suggests an unconventional origin for the quantum Hall state, a possibility we return to in light of additional experimental evidence discussed below.
Next, we investigate the unusual robustness of the superconducting phase against $B_{\perp}$ in Figs. 3a–b, which present transport measurements across $n$$B$ planes intersecting regime I at fixed displacement fields. A pronounced resistive peak in $R_{xx}$ appears within regime I around $B_{\perp}=1$ T, marking the lower-field boundary for both the superconducting and quantum Hall states. We interpret this feature as indicative of a transition in the underlying electronic order.
@@
On the high-density side of regime I, quantum Hall states emerge along well-defined linear trajectories in the $n$$B$ plane, marked by black dashed lines in Fig. 3a–b. These trajectories extrapolate back to the low-density boundary of regime I at $B_{\perp}=0$, forming a characteristic pattern reminiscent of a Landau fan diagram Zhang et al. (2005). Figure M4a presents a line trace cutting through this sequence of quantum Hall states. The evolution of $R_{xx}$ as a function of both $B_{\perp}$ and $n$ shows excellent agreement with Shubnikov–de Haas (SdH) oscillations, providing further confirmation of a robust quantum Hall sequence.
The slopes of the quantum Hall trajectories allow us to extract the Streda slopes, revealing an intriguing hierarchical structure: the sequence of quantum Hall states occupy both integer and half-integer Landau level (LL) fillings, $\nu=N/2$ with $N\in\mathbb{Z}$.
In conventional two-dimensional electron systems, incompressible states at half-filled LLs—commonly referred to as even-denominator fractional quantum Hall effects (FQHE)—have attracted significant interest due to their non-trivial topological nature. These states typically exhibit much smaller energy gaps than neighboring integer quantum Hall (IQHE) states, and their stability is strongly dependent on $B_{\perp}$.
Interestingly, the quantum Hall states within regime I do not exhibit the expected dependence on $B_{\perp}$. Instead, they show sharply defined temperature onsets that remain nearly constant over a broad range of $B_{\perp}$, as shown in Fig. 3c, indicating that their stability is largely independent of magnetic field. Moreover, the onset temperatures at both integer and half-integer filling factors are nearly identical, suggesting comparable robustness. These observations point to an unconventional hierarchical structure, which is further supported by the Shubnikov–de Haas (SdH) oscillations measured across the quantum Hall sequence (see Fig. M4). A Fourier analysis of the oscillations reveals a dominant spectral peak at $f_{\nu}=2$. Taken together, these findings indicate that, within regime I, the integer and half-integer states belong to a single, unified quantum Hall sequence.
@@
It is worth noting that the temperature dependence of the quantum Hall states closely mirrors that of the superconducting phase. As shown in Fig. 3d, the superconducting transition remains essentially unchanged across a broad range of $B_{\perp}$, and its transition temperature closely matches the onset temperature of the adjacent quantum Hall states.
Figure 4: Transition temperature and critical current. (a) Longitudinal resistance $R_{xx}$ (left panel) and transverse resistance $R_{xy}$ (right panel) measured as functions of carrier density $n$ and temperature $T$ across regime I at $B_{\perp}=3$ T and $D=1000$ mV/nm. (b) Differential resistance $dV/dI$ plotted as a function of $n$ and d.c. current bias $I_{\text{DC}}$ across regime I, measured at the same $B_{\perp}$ and $D$. Data in the left (right) panel were obtained in the longitudinal (transverse) measurement configuration.
@@
Figure 4a presents an $n$$T$ map of the transport response across regime I, measured at fixed $B_{\perp}$ and displacement field $D$. The superconducting transition and the onset of the quantum Hall effect together trace a continuous envelope that evolves smoothly with carrier density $n$. A similar correspondence is observed in the current–voltage (I–V) characteristics. As shown in Fig. 4b, the superconducting state exhibits a peak in the differential resistance $dV/dI$, which defines a critical supercurrent that varies with $n$ across regime I.
Remarkably, the quantum Hall state on the high-density side also displays strong sensitivity to d.c. current bias: a dip in $dV/dI$ marks a critical current that separates a highly nonlinear response at low bias from a linear response at high bias. This critical current in the quantum Hall regime closely tracks the evolution of the critical supercurrent in the adjacent superconducting phase. Together, these observations strongly support a direct connection between the mechanisms underlying the superconducting and quantum Hall phases within regime I.
Together, our findings present a compelling challenge to existing models of superconductivity and quantum Hall physics. Any successful theoretical framework must account for the following key observations:
@@
First, a superconducting phase is stabilized—rather than suppressed—by an out-of-plane magnetic field $B_{\perp}$. That the superconducting phase persists beyond $B_{\perp}>4$ T is unprecedented in two-dimensional superconductors. Second, a sequence of quantum Hall states emerges at both integer and half-integer filling factors, forming a single unified hierarchy. Third, the unique interplay between the superconducting and quantum Hall phases—including their shared onset temperature and adjacent locations in phase space—suggests a common underlying mechanism. Finally, the boundaries of regime I follow a Streda slope of $t=2.5$, consistent with a fractional quantum Hall state, yet without signatures of an incompressible state.
A possible interpretation of the Streda slope involves two components: a compressible state formed by carriers in a partially filled band, and a gapped, interaction-driven state characterized by a nontrivial Chern number of 2.52.52.5. In this picture, lines of constant Streda slope correspond to fixed charge density within the compressible band. The boundary of the cylindrical feature shown in Fig. 1b follows the same contour in the $D$–(charge density in the partially filled band) plane for each value of magnetic field.
Building on this framework, we now consider two candidate scenarios that may account for the observed superconducting and quantum Hall states.
The first scenario invokes the intriguing possibility of a quantum Hall effect of bosons. In this picture, the application of an out-of-plane magnetic field stabilizes pre-formed Cooper pairs—bound states of two electrons—within regime I. On the low-density side of regime I, a superconducting phase emerges as these bosonic pairs acquire long-range phase coherence. The remarkable stability of this superconducting phase against strong out-of-plane magnetic fields suggests that the Cooper pairs form in an odd orbital angular momentum channel, leading to an unconventional order parameter. On the high-density side, quantum Hall states arise when the bosonic Cooper pairs occupy their own Landau levels, in contrast to conventional models in which single electrons fill fermionic Landau levels. This bosonic quantum Hall scenario has been extensively explored in theoretical proposals  Cooper (2020) and realized in cold atom platforms  Léonard et al. (2023), but its manifestation in solid-state systems has remained elusive until now.
As shown in Fig. 3a–b, the boundary of the superconducting phase exhibits slight dips along the trajectories of the quantum Hall states, indicating that superconductivity is suppressed at the onset of quantum Hall states. This observation suggests that the superconducting and bosonic quantum Hall phases are competing orders within the same underlying bosonic framework.
This bosonic framework provides a natural explanation for the experimentally observed connection between superconducting and quantum Hall states: both emerge at similar magnetic fields, where Cooper pairs are stabilized, and both are suppressed at approximately the same temperature, $T\approx 0.5$ K, where Cooper pairs are destroyed. This also accounts for the $B$-independence of the onset temperature, as the Cooper pairing instability is not directly influenced by the magnetic field.
Moreover, the picture of competing bosonic phases offers a compelling explanation for the hierarchical structure of the quantum Hall sequence at both integer and half-integer filling factors—features that are less straightforward in a fermionic Landau level description. In fermionic systems, the energy scales for integer and fractional quantum Hall effects are governed by single-particle gaps and interaction energies, respectively, leading to marked differences in their robustness. In contrast, the quantum Hall effect in bosonic systems necessarily arises from interactions, implying that both integer and fractional bosonic states are governed by a common, interaction-driven energy scale. We note that filling factors discussed here are expressed in terms of electron density. Since Cooper pairs carry twice the charge of electrons, their effective filling factor is reduced by a factor of four.
In this picture, the Streda slope of $t=2.5$ observed in regime I reflects a topological gap located below the Fermi surface. While this topological structure is not directly responsible for the emergence of superconductivity and the quantum Hall effect, it modifies the charge carrier density and the dielectric environment in a manner that is favorable for stabilizing the Cooper pairing instability.
We also consider a second scenario in which a fractional Chern insulator (FCI) governs the physics of regime I. In this picture, regime I is populated by anyonic quasiparticles, as implied by the fractional Streda slope $t=2.5$. The quantum Hall states appearing at electronic filling $\nu=N/2$ is then interpreted as a sequence of quantum Hall states formed by these anyonic quasiparticles. In turn, superconductivity emerges as another correlated phase of anyons, leading to an anyonic superconductor—an exotic and theoretically rich phase that has recently attracted considerable interest Xu and et al. (2025); Nosov et al. (2025); Pichler et al. (2025). Within this framework, the characteristic onset temperature $T\approx 0.5~\textrm{K}$ reflects the energy scale necessary to stabilize the underlying FCI.
While this scenario appears consistent with most experimental observations, it faces several notable challenges. For example, effective Landau levels (LLs) of anyonic quasiparticles have not been identified to form such an extensive sequence, making their realization here highly nontrivial. Furthermore, filling LLs up to $\nu=6.5$ requires a substantial density of quasiparticles. Achieving such a high quasiparticle population while maintaining the robustness of a fractional Chern insulator (FCI) is highly unusual and presents a serious constraint for this scenario. In particular, this scenario raises a question regarding the Chern number of the underlying topological band.

### v2

@@ A Hierarchy of Topological and Superconducting States in Rhombohedral Hexalayer Graphene
Figure 1: Multi-dimensional phase space of rhombohedral hexalayer graphene. (a) Schematic of the R6G heterostructure. The red arrow indicates the out-of-plane magnetic field $B_{\perp}$. (b) Schematic phase diagram defined by carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. (c, d) nnDD maps of longitudinal resistance measured at $B=0$, with current flowing along the principal axes of (c) minimum and (d) maximum conductivity. (e) Annotated nnDD map highlighting prominent phases corresponding to panels (c, d). (f, g) nnBB maps measured at $D=967$ mV/nm, showing (f) longitudinal resistance $R_{xx}$ and (g) transverse resistance $R_{xy}$. White dashed lines trace the trajectories of integer quantum Hall states, exhibiting integer Streda slopes. The boundaries of regime I, highlighted by the black dashed lines, follow a Streda slope of $t=2.5$. In panel (f), white dotted lines show Streda slopes of 2 and 3, which clearly deviate from the boundaries of regime I.
@@
Figure 2: $B_{\perp}$-stabilized superconductivity and coexisting quantum Hall state. nnDD maps of (a) $R_{xx}$ and (b) $R_{xy}$ measured at different values of $B_{\perp}$. The dashed line outlines the boundary of regime I, while the dash-dotted line encloses the superconducting phase. (c) Line cuts of $|R_{xy}|$ as a function of $n$ at $D=1000$ mV/nm, measured at various values of $B_{\perp}$. (d, e) Temperature dependence of $R_{xx}$ (red) and $R_{xy}$ (blue) measured from (d) the $B_{\perp}$-stabilized superconducting phase and (e) the adjacent quantum Hall state.
@@
Figure 3: An exotic sequence of quantum Hall states. (a, b) nnBB map of $R_{xx}$ (top panel) and $R_{xy}$ (bottom panel) measured at (a) $D=980$ mV/nm and (b) $D=1000$ mV/nm. The superconducting region is enclosed by a white dashed line in the top panel, while the sequence of quantum Hall states forms a Landau fan, marked by black dashed lines.
An enduring theme in the study of superconductivity is its fundamental antagonism with magnetic fields: superconducting phases are typically suppressed by even modest applied fields. It is therefore striking when superconductivity persists under large magnetic fields, defying conventional wisdom. Such behavior sets high-$T_{c}$ superconductors apart from their more mundane counterparts, establishing a vibrant research frontier over the last decades  Phillips et al. (2021). Even more exotic is the scenario in which a superconducting phase is not merely resilient to magnetic fields, but is in fact stabilized by them—pointing to an unconventional pairing symmetry. A prominent example is UTe2, a candidate spin-triplet superconductor that exhibits field-induced reentrant superconductivity under strong magnetic fields Ran et al. (2019a, b); Aoki et al. (2019); Lewin et al. (2023).
@@
When superconductivity persists to high magnetic fields, it opens the door to a rare and intriguing regime in which superconductivity coexists with another hallmark of two-dimensional (2D) electron systems: the quantum Hall effect. The interplay between these two phenomena has long been proposed as a key platform for realizing fault-tolerant quantum computation, owing to the potential emergence of non-Abelian anyons and topological superconductivity Nayak et al. (2008); Das Sarma et al. (2015); Peralta Gavensky et al. (2020); Alicea and Fendley (2016). Over the past decade, this coexistence has been explored extensively in engineered hybrid systems, where superconducting and quantum Hall regions are coupled across artificial interfaces Amet et al. (2016); Lee et al. (2017); Gül et al. (2022); Zhao et al. (2023), as well as in 2D material heterostructures hosting superconductivity and topological order at zero magnetic field Stepanov et al. (2020); Choi et al. (2025). Despite these advances, achieving an intrinsic intertwinement of superconductivity and the quantum Hall effect within a single 2D electronic system—without relying on interface engineering—remains a central open challenge. A recent experimental report has identified such coexistence near zero magnetic field, sparking renewed interest in the possibility of anyonic superconductivity and its topological implications  Xu et al. (2025); Nosov et al. (2025); Pichler et al. (2025); Shi and Senthil (2024).
@@
Facilitated by a dual-encapsulated heterostructure, as illustrated in Fig. 1a, the low-temperature phase space of R6G is governed by three independently tunable parameters: carrier density $n$, displacement field $D$, and out-of-plane magnetic field $B_{\perp}$. Superconductivity and the sequence of quantum Hall states are both confined within a unique boundary, which encircles a tilted cylindrical volume within this nnDDBB space, as illustrated in Fig. 1b. For simplicity, we will refer to this portion of the phase space as regime I.
@@
We begin by defining the boundary of regime I through its distinct transport signatures. To this end, we perform measurements across multiple two-dimensional slices of the nnDDBB parameter space. In particular, we examine nnDD planes at five different values of out-of-plane magnetic field $B_{\perp}$ (blue dashed lines in Fig. 1b), as well as nnBB planes at three fixed values of displacement field $D$ (red dashed lines in Fig. 1b).
@@
At zero magnetic field ($B_{\perp}=0$), the cross section of regime I is bounded by the electronic smectic phase and the striped superconducting state, both previously identified through angle-resolved transport studies Morissette et al. (2025a). The transport response across this regime is characterized in Figs. 1c–d, which show longitudinal resistance measured across the nnDD plane at $B_{\perp}=0$, along two orthogonal principal axes corresponding to maximum and minimum conductivity. The smectic phase occupies the region of pronounced transport anisotropy (see Fig. M2b), highlighted by the red-shaded area in Fig. 1e. Within this anisotropic regime, two superconducting phases—denoted SC i and SC ii—emerge concurrently with the smectic order and inherit its strong directional character  Morissette et al. (2025a).
@@
The $B_{\perp}=0$ cross-section of regime I appears as a metallic phase situated between SC i and SC ii, enclosed by black dashed lines in Figs. 1c–e. As $B_{\perp}$ increases, the boundaries of regime I trace a well-defined trajectory in the nnBB plane, marked by peaks in the longitudinal resistance $R_{xx}$ and abrupt jumps in the Hall resistance $R_{xy}$. These boundaries, indicated by black dashed lines in Fig. 1f-g, exhibit a Streda slope of $t=2.5$  Středa and Smrčka (1983); Widom (1983); MacDonald and Středa (1984).
@@
Within the stripe regime, two additional metallic phases—regime II and regime III—are highlighted in Figs. 1c–e. These phases resemble regime I in their transport characteristics and exhibit well-defined boundaries in the nnBB plane, following distinct Streda slopes: $t=2$ for regime II and $t=3.5$ for regime III (see Fig. M8 for further discussion). The presence of such sharply delineated, evolving boundaries is intriguing. Although a well-defined Streda slope often signals a topological phase, we do not observe transport features typically associated with incompressible quantum Hall states Huang et al. (2025). We revisit this issue later in the context of additional experimental evidence.
@@
Figures 2a–b show nnDD maps of the longitudinal and Hall resistances measured at various values of $B_{\perp}$. At each field, regime I is delineated by peaks in $R_{xx}$, defining a boundary that shifts systematically with $B_{\perp}$ in both carrier density $n$ and displacement field $D$. These evolving boundaries are marked by dashed lines in Figs. 2a–b. Most notably, upon applying a non-zero $B_{\perp}$, a superconducting pocket emerges on the low-density side of regime I, while a sequence of quantum Hall states appears on the high-density side.
The superconducting phase—enclosed by dash-dotted lines—is identified by vanishing longitudinal resistance $R_{xx}$ and zero Hall resistance $R_{xy}$, appearing as dark blue and white, respectively, in the color maps. Figure 2c presents line cuts of the transport response as a function of $n$, taken at fixed $D=1000$ mV/nm and varying $B_{\perp}$. At $B_{\perp}=0$, the system remains resistive across the entire density range. In contrast, at $B_{\perp}=2.5$ and $3.3$ T, a superconducting dome is clearly visible, characterized by a sharp drop to zero resistance. Upon further increasing $B_{\perp}$ to $5$ T, the superconducting phase is fully suppressed.
@@
Remarkably, both the superconducting and quantum Hall phases exhibit sharply defined onsets as a function of temperature (Fig. 2d–e). The RRTT curves in Fig. 2d display behavior consistent with a superconducting transition, characterized by a critical temperature $T_{c}\approx 0.5$ K. In contrast, the similarly sharp onset observed for the quantum Hall state in Fig. 2e is highly unusual. Conventional quantum Hall states typically display a gradual thermal activation in $R_{xx}$, while $R_{xy}$ remains largely quantized and temperature-independent (see Fig. M6). The abrupt temperature dependence observed here suggests an unconventional origin for the quantum Hall state, a possibility we return to in light of additional experimental evidence discussed below.
Next, we investigate the unusual robustness of the superconducting phase against $B_{\perp}$ in Figs. 3a–b, which present transport measurements across nnBB planes intersecting regime I at fixed displacement fields. A pronounced resistive peak in $R_{xx}$ appears within regime I around $B_{\perp}=1$ T, marking the lower-field boundary for both the superconducting and quantum Hall states. We interpret this feature as indicative of a transition in the underlying electronic order.
@@
On the high-density side of regime I, quantum Hall states emerge along well-defined linear trajectories in the $n\text{--}B$ plane, tracing minima in $R_{xx}$ and maxima in $|R_{xy}|$. These trajectories, marked by black dashed lines in Fig. 3a–b, form a characteristic pattern reminiscent of a Landau fan diagram Zhang et al. (2005). However, the sequence of quantum Hall states in this regime exhibits several unusual features that set it apart from conventional Landau fans.
In particular, the observed trajectories follow Streda slopes corresponding to both integer and half-integer Landau level (LL) fillings, with $\nu=N/2$, where $N\in\mathbb{Z}$. This pattern would conventionally imply the emergence of incompressible states at fully and half-filled LLs. However, a striking discrepancy arises between the inferred filling factors and the actual Hall quantization. As shown in Fig. 3c–e, line cuts along representative trajectories reveal vanishing $R_{xx}$, signaling fully developed incompressible states, while the Hall resistance $R_{xy}$ remains quantized at integer multiples of the conductance quantum.
This mismatch between the Streda slope and Hall quantization suggests an unconventional origin. One possibility is a reduction in the effective LL degeneracy by a factor of two, such that a sequence of integer quantum Hall states occurs along Streda slope at $\nu=N/2$.
In addition, the observed quantum Hall states exhibit a highly unusual hierarchical structure. In conventional two-dimensional electron systems, incompressible states at half-filled Landau levels—commonly associated with even-denominator fractional quantum Hall effects (FQHE)—possess significantly smaller energy gaps than their integer counterparts and exhibit strong dependence on the out-of-plane magnetic field $B_{\perp}$. By contrast, quantum Hall states in regime I display comparable robustness between integer and half-integer Streda slopes, evidenced by nearly identical onset temperature, as illustrated in Fig. 4a. This offers further indication that the underlying incompressible states are IQHE in nature.
@@
Moreover, the onset temperature of quantum Hall states remain nearly constant across a wide range of $B_{\perp}$. This temperature dependence closely mirrors that of the superconducting phase. As shown in Fig. 4b, the superconducting transition remains essentially unchanged across a broad range of $B_{\perp}$, and its transition temperature closely matches the onset temperature of the adjacent quantum Hall states.
Figure 4: Transition temperature and critical current. (a, b) Temperature dependence of $R_{xy}$ measured at different $B_{\perp}$ values within (a) quantum Hall states and (b) the superconducting phase. (c) Longitudinal resistance $R_{xx}$ (left panel) and transverse resistance $R_{xy}$ (right panel) as functions of carrier density $n$ and temperature $T$ across the superconducting and quantum Hall states at $B_{\perp}=3$ T and $D=1000$ mV/nm. (d) Differential resistance $dV/dI$ as a function of $n$ and d.c. current bias $I_{\mathrm{DC}}$ at $B_{\perp}=3$ T and $D=1000$ mV/nm. Data in the left (right) panel were obtained in the longitudinal (transverse) measurement configuration. (e–g) Schematic diagrams illustrating possible mechanisms underlying superconductivity and quantum Hall states. (e) Regime I is associated with the emergence of a CDW order, with insulating islands shown as white circles. (f, g) Away from the low-density boundary of regime I, itinerant electrons occupy the regions between insulating islands, giving rise to (f) superconductivity on the low-density side of regime I and (g) the quantum Hall effect on the high-density side of regime I.
@@
A close correlation between the quantum Hall and superconducting states is further illustrated in Fig. 4c–d. An $n\text{--}T$ map of the transport response, taken at fixed $B_{\perp}$ and displacement field $D$, reveals that the superconducting transition and the onset of the quantum Hall effect trace a continuous envelope that evolves smoothly with carrier density $n$. A similar correspondence is observed in the current–voltage (I–V) characteristics, as shown in Fig. 4d. The quantum Hall state, which appears on the high-density side of regime I, exhibits pronounced nonlinearity in its transport response, evidenced by large variations in the differential resistance $dV/dI$ in both longitudinal and transverse channels. This nonlinearity defines a sharp threshold in the d.c. current bias, which closely aligns with the critical current of the superconducting state—marked by a peak in $dV/dI$. Taken together, these coordinated behaviors strongly suggest a shared origin underlying both the superconducting and quantum Hall states.
@@
In the following, we discuss potential mechanisms that may underlie the intriguing coexistence of superconductivity and quantum Hall states. Any viable explanation must account for the following key experimental observations. First, a superconducting phase is stabilized—rather than suppressed—by an out-of-plane magnetic field $B_{\perp}$. That the superconducting phase persists beyond $B_{\perp}>4$ T is unprecedented in two-dimensional superconductors. Second, the sequence of quantum Hall states emerges at both integer and half-integer filling factors, displaying a prominent mismatch between the Streda slope and Hall quantization. Third, the unique interplay between the superconducting and quantum Hall phases—including their shared onset temperature and adjacent locations in phase space—suggests a common underlying mechanism.
We consider a scenario involving charge density wave (CDW) order, wherein electrons are localized to form an array of insulating islands, schematically illustrated as white circles in Fig. 4e–g. In this picture, the extent of regime I is governed by the stability of the CDW order across the $n\text{--}D\text{--}B$ phase space. As the carrier density $n$ increases from the low-density boundary of regime I, itinerant electrons are introduced occupying the regions between the insulating islands. These itinerant electrons contribute to conductive transport at $B_{\perp}=0$ and enable the emergence of superconductivity and quantum Hall states at $B_{\perp}>1$ T. The boundary of the CDW phase traces a finite Streda slope in the $n\text{--}B$ plane, suggesting that the reconstructed electronic bands feature non-zero Berry curvature. However, due to the mismatch between Streda slope and Hall plateau, the exact Streda slope likely is not representing the actual Berry curvature.
In the presence of an out-of-plane magnetic field, LL formation of the itinerant electrons gives rise to the sequence of quantum Hall states observed in regime I (Fig. 4f). Within this framework, these quantum Hall states resemble a re-entrant integer quantum Hall effect Eisenstein et al. (2002); Xia et al. (2004); Halperin and Jain (2020). Because electrons localized within the insulating islands do not contribute to transport, the effective carrier density is reduced. This reduction yields Hall plateaus comparable to those of the IQHE emanating from the CNP, indicating the same effective density occurring at distinct values of the actual carrier density (Fig. 2b and Fig. 3a–b).
Crucially, the emergence of CDW order decouples the Streda slope from the quantized Hall plateau. When extra charge carriers are added to the sample, they preferentially occupy the regions between insulating islands. As a result, the effective density of itinerant electrons, $n^{\ast}$, increases at a faster rate compared to the actual electron density $n$. This naturally accounts for the mismatch between the Streda slope and the quantized Hall plateau.
This framework leads to an intriguing scenario in which superconductivity coexists with charge density wave (CDW) order, occupying the regions between insulating CDW islands (Fig. 4g). These metallic regions supports superconducting transport at low temperatures. The close proximity of the superconducting and quantum Hall phases in the phase diagram—and their nearly identical onset temperatures—suggests that both originate from the same population of itinerant electrons, with superconductivity (quantum Hall states) stabilized in the low-density (high-density) limit. As the temperature increases, the CDW order melts, disrupting the spatial segregation between localized and itinerant carriers. This melting transition naturally accounts for the sharply defined thermal onset observed in both superconducting and quantum Hall responses (Fig. 4a–b).